Write out the estimated regression equation.

Test for the significance of the slope at  LaTeX: \alpha\:=.05

Determine the coefficient of correlation between y and y.

For questions 26-28 use the following information: Below you are given a partial computer output based on a sample of 21 observations relating an independent variable (x) and a dependent variable (y). Predictor Coefficient Standard Error Constant 30.139 1.181 0.252 0.022 ANOVA Sum of Squares Source 1759.481 Model 259.186 Error

y and y.
Answer #1

Estimated regression equation:

y = 30.139 - 0.252 x

--------------------------------------

Slope, b = -0.252

Standard error, se(b1) = 0.022

Null and alternative hypothesis:

Ho:β₁ = 0; H1:β₁ ≠ 0

n = 21

df = n-2 = 19

α = 0.05

Test statistic:

t = b /se(b1) = -0.252/0.022 = -11.4545

p-value = T.DIST.2T(-11.4545, 19) = 0.0000

Decision : p-value < α Reject the null hypothesis.

--------------------------------------

Coefficient of correlation, R2 =

1759.481 R2SSR = 0.8716 SSR+SSE 1759.481259.186